OpenStudy (xxstudyingstuffxx):

What is the relationship between 9.125 • 10^-3 and 9.125 • 10^2. (Simpler equation in comments.) Fan + Medal

OpenStudy (xxstudyingstuffxx):

What is the relationship between \[9.125 \times 10^{-3} \] and \[9.125 \times 10^{2}\]

OpenStudy (xxstudyingstuffxx):

@AloneS @518nad @zyi6

OpenStudy (xxstudyingstuffxx):

Anyone?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (xxstudyingstuffxx):

Uh, not really. Are they asking what is the same about the two problems?

OpenStudy (skullpatrol):

Yes they are asking that too. Recall $$\huge a^m • a^n = a^{m+n}$$

OpenStudy (skullpatrol):

Replace a with 10.

OpenStudy (xxstudyingstuffxx):

I found it out. Thankyou

OpenStudy (skullpatrol):

Thanks for trying to learn :-)